shift summary, evening 5 oct 99

Marj Corcoran (corcoran@physics.rice.edu)
Wed, 6 Oct 1999 00:04:02 -0500 (CDT)


Shift Summary for Evening shift, 5 Oct 99

On Shift: Leo until 9:30PM, Marj from 7:30PM on

Hours of beam available at all intensities: 0
Hours of beam available at nominal intensity: 0
Hours of E799 data (special runs notv included): 0
Hours of special runs (laser scans, muon runs...): 0

Accounting of lost time: Beam off due to CHL problems

========================================================================

Runs ended this shift:
Run 14803--cosmics, not to tape

=======================================================================

Tapes brought to FCC this shift: None

==================================================================

Description of shift: No beam. Leo measures frequency of beam
modulation, see below.

===================================================================

16:00 DN fills alcohol bubbler, level is now 3.5"

20:00 Marj lowers chamber voltages to 2000V for all planes, all views,
as well as the window wires. The sdaq thinks any voltage below
2000V is "off", so, in order for the window wires supply to not appear
red on the sdaq screen, the window wires were also set to 2000V.
The chamber plots on the display will now look bad, don't worry
about it. The sdaq still occasionally thinks a chamber is off, but it
goes away in a few seconds.

21:00 Marj adds kumac to da_ktev kumac area. dc1cur#live [run]
It plots live time, boa, sem and DC1 current vs spill (as in
l3#live), then projects DC1 current and fits a Gaussian. Try it.

21:30 Leo succeeds in observing 2Hz structure in beam by doing FFT
of the structure in MU2 from a muon run. The exact period turns
out to be 2.215 Hz. There seems also to be a beat freqency indicating
two frequencies separated by 90mHz. See Log XXII, pg 62-64. Way cool.

22:30 Gas check list done, what there is of regular check list done.
Note that the three Ar/ethane spares are the unanalyzed grade.
Has more been ordered?

24:00 Rick F arrives, had off to owl.
=======================================================================

Items to follow up on next shift: New copier did not arrive on this shift.
Owl shift should keep a sharp eye out for the new copier.

There is a large quantity of lasagania in the fidge, owl shift/day shift
should consume it. Thanks to Leo (whose last name ends in a vowel).

=======================================================================